OpenStudy (anonymous):

help please. use the formula for the area of a trapezoid. show work how many square feet of grass are there on a trapezoidal field with a height of 75 ft and bases of 124 ft and 81 ft?

OpenStudy (igreen):

Do you know what the formula is? @allen10

OpenStudy (k_lynn):

This file has came in handy for me this past year. You might want so save it for future reference. Just plug in the numbers and solve

OpenStudy (igreen):

\(Area = \dfrac{(b_1 + b_2) \times h}{2}\)

OpenStudy (igreen):

\(Area = \dfrac{(b_1 + b_2) \times h}{2}\) We are given the height and the two bases. So input them into the formula: \(Area = \dfrac{(124 + 81) \times 75}{2}\) Add: \(Area = \dfrac{205 \times 75}{2}\) Multiply: \(Area = \dfrac{15,375}{2}\) Divide: \(Area = 7687.5\)

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @allen10